# Global Climate Model Bias Correction Using Deep Learning

### v1

@@ 5.2 Comparison of models in future projections
Figure 5: Monthly sea surface temperature (SST) in the BoB during 2021 Dec to May, comparing ORAS5 reanalysis data, raw CNRM-CM6 model projections (CNRM-CM6), UNet-corrected SST (UNet), EDCDF corrected SST (EDCDF), BiLSTM corrected SST (BiLSTM), ConvLSTM corrected SST (ConvLSTM).
@@
Figure 6: Monthly sea surface temperature (SST) in the BoB during 2021 June to Nov, comparing ORAS5 reanalysis data, raw CNRM-CM6 model projections (CNRM-CM6), UNet-corrected SST (UNet), EDCDF corrected SST (EDCDF), BiLSTM corrected SST (BiLSTM), ConvLSTM corrected SST (ConvLSTM).
@@ 6 Conclusion
As a baseline, linear regression models and statistical EDCDF methods were also evaluated. We found that baseline models provide basic correction capabilities but fail to capture the complex nonlinear relationships present in climate projections. Although ConvLSTM produced reasonable annual mean projections aligned with statistical methods, only the U-Net model consistently captured both spatial and temporal variability, outperforming all alternatives in reducing errors in the climate projections. This confirmed U-Net as the most suitable architecture for GCM error correction in the Bay of Bengal. Performance evaluation demonstrated a 15% reduction in root mean square error (RMSE) in SST projections compared to the widely used Equidistant Cumulative Distribution Function (EDCDF) method. This improvement highlights the ability of deep learning approaches to model complex, nonlinear, and spatially dependent bias structures inherent in global climate models.
@@
All the data used in this work is available for download from the respective data providers with links provided in the manuscript. Our code and data produced from this work are available on GitHub [^2] https://github.com/AbhiPasula/Climate-Model-SST-Bias-Correction-Toolkit.
@@ Acknowledgment
We thank the Ministry of Earth Sciences and the Anusandhan National Research Foundation, Government of India, for their grants that allowed the purchase of the compute resources used in model training. The authors thank the members of the QUEST Lab at IISc for insightful discussion.

### v2

@@ 5.2 Comparison of models in future projections
Figure 5: Monthly sea surface temperature (SST) in the Bay of Bengal during 2021 Dec to May, comparing ORAS5 reanalysis data, raw CNRM-CM6 model projections (CNRM-CM6), UNet-corrected SST (UNet), EDCDF corrected SST (EDCDF), BiLSTM corrected SST (BiLSTM), ConvLSTM corrected SST (ConvLSTM).
@@
Figure 6: Monthly sea surface temperature (SST) in the Bay of Bengal during 2021 June to Nov, comparing ORAS5 reanalysis data, raw CNRM-CM6 model projections (CNRM-CM6), UNet-corrected SST (UNet), EDCDF corrected SST (EDCDF), BiLSTM corrected SST (BiLSTM), ConvLSTM corrected SST (ConvLSTM).
@@
5.4 Limitations
The present UNet model (and other architectures) work well with the particular climate model used here, namely, CNRM-CM6. However, when another climate model output is used for bias correction with the model trained on CNRM-CM6, the performance is not as good as a dedicated model trained for that climate model. A study using UNet trained with the CNRM-CM6 forecast, when applied with HadGEM3 (Andrews et al., 2020), shows few discrepancies compared to ORAS5, particularly in the transition months between seasons and extreme warming. A possibility for future research is to attempt fine-tuning or transfer learning from a bias corrector trained on one climate model and use it for another. A different approach is to employ multiple climate models during the training itself to leverage the generalizability of neural operators.
@@ 6 Conclusion
As a baseline, linear regression models and statistical EDCDF methods were also evaluated. We found that baseline models provide basic correction capabilities but fail to capture the complex nonlinear relationships present in climate projections. Although ConvLSTM produced reasonable annual mean projections aligned with statistical methods, only the U-Net model consistently captured both spatial and temporal variability, outperforming all alternatives in reducing errors in the climate projections. This confirmed U-Net as the most suitable architecture for GCM error correction in the Bay of Bengal. Performance evaluation demonstrated a 15% reduction in RMSE and a 12% increase in PCC in SST projections compared to the widely used Equidistant Cumulative Distribution Function (EDCDF) method. This improvement highlights the ability of deep learning approaches to model complex, nonlinear, and spatially dependent bias structures inherent in global climate models.
@@
All data used in this work are available for download from the respective data providers, with links provided in the manuscript. Our code and data produced from this work are available on GitHub (https://github.com/AbhiPasula/Climate-Model-SST-Bias-Correction-Toolkit).
@@ Acknowledgment
The authors thank the Ministry of Earth Sciences for the grant that allowed the purchase of the compute resources used in model training. The authors thank the members of the QUEST Lab at IISc for insightful discussion.
Statements and Declarations
Funding Information
This work was supported by the Ministry of Earth Sciences (grant number MoES/36/OOIS/Extra/84/2022).
Competing Interests
The authors have no relevant financial or non-financial interests to disclose.